artB What Celsius temperature, T, is required to change the volume of the gas sample in Part A (Ti 12 °C. V 1.77 103 L ) to a volume of 3.54x103 L? Assume no change in pressure or the amount of gas in the balloon Express your answer with the appropriate units. Hints Th=11 Value Units Submit My Answers Give Up
Answer #1

Ans. Charles’ Law: Volume (V) of a gas in a closed vessel is directly proportional to absolute temperature (T) at constant pressure. V/T for a gas is a constant at specified pressure.

                                    (V1/ T1) = (V2/ T2)            - equation 1

Where, V1 = 1.77 x103 L at T1=120C = 285.15 K

V2 = 3.54 x103 L at T2

Putting the values in equation 1-

            (1.77 x 103 L) / 285.15 K = (3.54 x 103 L) / T2

            Or, T2 = (3.54 x 103 L) / [(1.77 x 103 L) / 285.15 K] = 570.3 K

Therefore, T2 = 570.3 K = 297.150C = 2970C